what is the answer to: y=negative x +2 and 2y=4-2x
Found 3 solutions by hokies, richwmiller, Alan3354:
Answer by hokies(65)

The two equations are not independent from each other (the second equation is two times the first equation), so there are an infinite number of solutions to this system of equations.

Answer by richwmiller(17219)

Depends! What is the question?
Do you want a solution to the two equations where they intersect?
Do you want to know if they intersect?
Do you want their slopes?
Do you want their intercepts?
It would also help if you wrote them so they could be read. Why did you write out negitive (sic)?
sic means I copied the word as it was written and I know it is wrong.
Which part is negative? (x) or (x+2)
y=-x+2
2y=4-2x
for now I will assume just x

multiply equation 1 by 2
2y=-2x+4
rearrange
2y=4-2x
They are equations for the same line and do not intersect.
Therefore there is not enough information to solve them.
They both have the same set of solutions.

Answer by Alan3354(69443)

Parentheses are free. Use some.